OpenStudy (anonymous):

Does anyone know why this is true? The sum from n = -infinity to infinity of delta(t-n)/2? Not sure if LaTeX works on this so I'll write it out below.

OpenStudy (anonymous):

\[f(t) = \sum_{n =-\infty}^{\infty} \Delta(t-n)/2\]

OpenStudy (anonymous):

Sorry here is the equation I need explained: \[f(t) = \sum_{n = -\infty}^{\infty} (t-n)/2 = 1\]

OpenStudy (anonymous):

Sorry again :/ \[f(t) =\sum_{n = -\infty}^{\infty} \Delta(t-n)/2 = 1\]

OpenStudy (experimentx):

what's that delta

OpenStudy (anonymous):

I think it just means "change in", like a differential. So my interpretation of this is the infinite sum of changes in the function (t-n)/2... I just don't understand how that all adds up to one though.